OpenStudy (kmullis6):

Suppose M is between L and N. Use the Segment Addition Postulate to solve for the variable. Then find the lengths of LM, MN, and LN.

OpenStudy (kmullis6):

LM = 3x + 8 MN = 2x - 5 LN = 23

OpenStudy (dinamix):

x=4 ,LM+MN = LN

OpenStudy (kmullis6):

In the back of the book, the answer says 20, 3, 23.

OpenStudy (dinamix):

yup this i find it

OpenStudy (dinamix):

\[3x+8 +2x-5 = 23\]

OpenStudy (dinamix):

\[5x+3=23\]

OpenStudy (dinamix):

\[5x=23-3 =20 , x= 20/5 , x=4 \]

OpenStudy (dinamix):

LM = (3*4)+8 =20

OpenStudy (dinamix):

@kmullis6 did u see my steps ?

OpenStudy (kmullis6):

Yea I get it. Thank you for the help!

OpenStudy (dinamix):

np @ u are welcome thats make me feel good , so did u understand it or no @kmullis6

OpenStudy (kmullis6):

I understood what you said. If I need help with any of the others, is it ok to ask you again for help? @dinamix

OpenStudy (dinamix):

ok "_" @kmullis6